# Martin’s Conjecture in the Enumeration Degrees

### v1

@@ 6. A Non-Uniformly Invariant Function
A point of interest around Martin’s Conjecture is the dividing line between uniformly and non-uniformly $T$-invariant functions (on a cone). Early on, Steel conjectured that, under $AD$, they are exactly the same [27]. Steel’s conjecture remains open and a positive result implies Martin’s Conjecture. An unexpected consequence of Theorem 5.1 is that Steel’s conjecture fails for $e$-invariant functions. Notably, the example that we will construct works in $ZF$ and comes from Kalimullin pairs, a technical tool generalizing semicomputable sets that has been repeatedly used to obtain definability results in the enumeration degrees [24, 6, 13].
Semicomputable sets were introduced by Jockusch [15] to study the interaction between strong reducibilities. A set $A$ is semicomputable if $A$ is a left cut of a computable linear order. Note that if $A$ is semicomputable, $\overline{A}$ also is [^1] If $A$ is a left cut of a computable linear order, then $\overline{A}$ is a left cut of the inverse order.. We say that $\{A,\overline{A}\}$ form a semicomputable pair. Jockusch [15] proved that every Turing degree contains a nontrivial semicomputable pair; that is, one where $A$ and $\overline{A}$ are not c.e. In the enumeration degrees, the halves of nontrivial semicomputable pairs have different degrees and every total degree is the join of the degrees of the two halves of a semicomputable pair. Moreover, if $A$ is semicomputable then its enumeration degree is quasiminimal—the only total degree it bounds is $\bm{0}_{e}$—and forms a strong type of minimal pair with $\overline{A}$: the degree of any set $X$ is the greatest lower bound of the degrees of $A\oplus X$ and $\overline{A}\oplus X$ [3]. Kalimullin then introduced the $\mathcal{K}$-pairs (under the name of $e$-ideal pairs, for reasons that will become apparent below) to prove that the enumeration jump is definable [24]. They generalize the properties of semicomputable pairs in the context of enumeration degrees.
@@
We say that the $\mathcal{K}_{U}$-pair is nontrivial if $A\nleq_{e}U$ and $B\nleq_{e}U$. If $U$ can be chosen to be $\varnothing$, we omit the subscript and say that $\{A,B\}$ is a $\mathcal{K}$-pair.
Nontrivial $\mathcal{K}_{U}$-pairs have many useful properties, the following proposition summarizes the main ones:
Proposition 6.2.
@@
    (Kalimullin [24]) If $C\leq_{e}B$, then $\{A,C\}$ is a $\mathcal{K}_{U}$-pair.
@@
    (Kalimullin [24]) If $\{A,C\}$ is a $\mathcal{K}_{U}$-pair, then $\{A,B\oplus C\}$ is a $\mathcal{K}_{U}$-pair.
  3. (c)
    (Kalimullin [24]) For all $X$, we have that $U\oplus X$ is a greatest lower bound of $\{A\oplus U\oplus X,B\oplus U\oplus X\}$.
  4. (d)
    (Andrews et. al. [2]) $A\leq_{e}B^{\diamond}\oplus U$ and $A^{\diamond}\leq_{e}B\oplus U^{\diamond}$
  5. (e)
    (Kalimullin [24]) The set $A\oplus U$ is a strong quasiminimal cover of $U$.
Observe that left and right halves of a nontrivial $\mathcal{K}_{U}$-pair are symmetric, so the proposition above holds equally if we interchange the roles of $A$ and $B$. The first item implies that the notion of $\mathcal{K}_{U}$-pairs is degree-theoretic, and the first and second combined state that set of $\mathcal{K}_{U}$-partners of half of a $\mathcal{K}_{U}$-pair forms an ideal in the enumeration degrees. The converse of the third item also holds, which gives the definability of $\mathcal{K}$-pairs in the enumeration degrees. Moreover, it shows that a $K_{U}$-pair $\{A,B\}$ is also a $K_{V}$-pair for any $V\geq_{e}U$ and that it is nontrivial if we also have that $V\ngeq_{e}A,B$.
A semicomputable pair $\{A,\overline{A}\}$ is not only a $\mathcal{K}$-pair, but a maximal $\mathcal{K}$-pair [12]: if $C\geq_{e}A$, $D\geq_{e}\overline{A}$, and $\{C,D\}$ is a $\mathcal{K}$-pair, then $A\equiv_{e}C$ and $\overline{A}\equiv_{e}D$. Cai et. al. [6] proved that these are the only maximal $\mathcal{K}$-pairs (up to enumeration degree), concluding that total degrees are definable. This can be relativized with a total oracle $U$ to obtain that the degrees of maximal $\mathcal{K}_{U}$-pairs are exactly the degrees of semicomputable pairs relative to $U$.
Now we have all the ingredients to construct the $e$-invariant function that is not uniformly invariant. The next lemma provides the combinatorial core of the counterexample.
@@
If $\bm{a}\leq\bm{b}$, then $\bm{u}=(\bm{a}\vee\bm{u})\wedge(\bm{b}\vee\bm{u})=\bm{a}\vee\bm{u}$. This means that $\bm{a}\leq\bm{u}$, so $\{A,B\}$ is a trivial $\mathcal{K}_{U}$-pair, contrary to our assumption. If instead $\bm{a}^{\diamond}\leq\bm{b}$, the monotonicity of the skip implies that $\bm{u}^{\diamond}\leq\bm{a}^{\diamond}$. Combined with Proposition 6.2.d,
@@
Now that we have a tool to prove that a function is not uniformly $e$-invariant, we devote the rest of the section to the construction of a counterexample that works on every cone. The main obstacle is finding, on input $A$, if there are sets $B$ and $U$ such that $\{A,B\}$ is a nontrivial $\mathcal{K}_{U}$-pair so that we can map $A$ to $B$ while at the same time ensuring that if $C\equiv_{e}A$, we make a consistent choice $D\equiv_{e}B$ as the output on $C$. The second part can be managed by restricting our attention to maximal $\mathcal{K}_{U}$-partners, as those inhabit a single degree. While we can only ensure that maximal $\mathcal{K}$-pairs exist when the base $U$ is total, that is enough for our purposes. The next lemma takes care of the first part.
@@
There is a Borel $e$-invariant function $F\colon\mathcal{P}(\mathbb{N})\to\mathcal{P}(\mathbb{N})$ that is not uniformly $e$-invariant.
@@
Of course, we need to ensure that it is single-valued and Borel. We will construct the function $F$ to follow, on input $A$, the following instructions. First, search for a total set $X\oplus\overline{X}$ below $A$ and a candidate $B$ below $A^{\diamond}$ such that $\{A,B\}$ is a maximal $K_{X\oplus\overline{X}}$-pair. If found, map $A$ to $B$. Otherwise, map $A$ to $A$. While the procedure to compute $F(A)$ from $A$ is not computable, it is arithmetical. To make things precise, we define $F(A)$ by the following procedure:
@@
A rough calculation using Theorem 2.6(III) in [24] gives a bound of $\Sigma^{0}_{4}(A^{\prime})$ on the complexity of $F(A)$. The uniqueness (up to $e$-degree) of maximal $\mathcal{K}$-partners guarantees that the function is $e$-invariant, and step 2 ensures that $F$ is not locally constant. Finally, Lemma 6.3 shows that $F$ is not uniformly $e$-invariant. ∎
@@ 7. Further Questions
Is there an $e$-invariant function $f$ such that $f$ is regressive and $f\restriction\deg(A)$} is not constant for all $A\subseteq\omega$?
The failure of the Cone Theorem in the enumeration degrees leaves open whether some of the consequences of determinacy can still be proven. For example, a classical application of TD shows that there is a cone $\mathcal{C}$ of Turing degrees such that every cone $\mathcal{D}_{T}(\geq\bm{d})$ above an element $\bm{d}\in\mathcal{C}$ has the same first-order theory. The same argument can be used to show that, under TD, there is a total enumeration degree $\bm{t}$ such that for all total degrees $\bm{d}\geq\bm{t}$, the cone of enumeration degrees $\mathcal{D}_{e}(\geq\bm{d})$ has the same first-order theory as $\mathcal{D}_{e}(\geq\bm{t})$. However, this is false for arbitrary enumeration degrees. For example, the proof of downwards density of $\mathcal{D}_{e}$ relativizes to any total degree; thus, for a total degree $\bm{t}$, $\mathcal{D}_{e}(\geq\bm{t})$ has no minimal degrees [14]. In contrast, by a theorem of Calhoun and Slaman [7], above every total degree there is a degree $\bm{a}$ such that $\mathcal{D}_{e}(\geq\bm{a})$ has a minimal degree.
@@
The same question might be asked about other classes of degrees. For example, the continuous degrees are strictly between the total and the cototal degrees under the subset relation and are first-order definable [1]. Given that all the proofs that there is a non-total continuous degree are topological in nature, understanding the theory of the cone above a non-total continuous degree could provide a useful technical tool.
@@ Acknowledgments
Thanks to Mariya Soskova for many helpful discussions and advice, to Josiah Jacobsen-Grocott for sharing his insight, and to Beth Gatho for her thoughtful suggestions. The author was supported by the Secretaría de Ciencia, Humanidades, Tecnología e Innovación of México through scholarship number 836694.

### v2

@@ 6. A Non-Uniformly Invariant Function
A point of interest around Martin’s Conjecture is the dividing line between uniformly and non-uniformly $T$-invariant functions (on a cone). Early on, Steel conjectured that, under $AD$, they are exactly the same [27]. Steel’s conjecture remains open and a positive result implies Martin’s Conjecture. An unexpected consequence of Theorem 5.1 is that Steel’s conjecture fails for $e$-invariant functions. Notably, we will construct a Borel example using the notion of Kalimullin pair, a technical tool generalizing semicomputable sets that has been repeatedly used to obtain definability results in the enumeration degrees [24, 6, 13].
6.1. Semicomputable sets and $\mathcal{{K}}$-pairs
Semicomputable sets were introduced by Jockusch [15] to study the interaction between strong reducibilities. A set $A$ is semicomputable if $A$ is a left cut of a computable linear order. Note that if $A$ is semicomputable, $\overline{A}$ also is [^1] If $A$ is a left cut of a computable linear order, then $\overline{A}$ is a left cut of the inverse order.. We say that $\{A,\overline{A}\}$ form a semicomputable pair. Jockusch [15] proved that every Turing degree contains a nontrivial semicomputable pair; that is, one where $A$ and $\overline{A}$ are not c.e. In the enumeration degrees, the halves of nontrivial semicomputable pairs have different degrees and every total degree is the join of the degrees of the two halves of a semicomputable pair. Moreover, if $A$ is semicomputable then its enumeration degree is quasiminimal—the only total degree it bounds is $\bm{0}_{e}$—and forms a strong type of minimal pair with $\overline{A}$: the degree of any set $X$ is the greatest lower bound of the degrees of $A\oplus X$ and $\overline{A}\oplus X$ [3]. Kalimullin then introduced the $\mathcal{{K}}_{U}$-pairs (under the name of $U$-$e$-ideal pairs, for reasons that will become apparent below) to prove that the enumeration jump is definable [24]. They generalize the properties of semicomputable pairs in the context of enumeration degrees.
@@ 6.1. Semicomputable sets and $\mathcal{{K}}$-pairs
We say that the $\mathcal{{K}}_{U}$-pair is nontrivial if $A\nleq_{e}U$ and $B\nleq_{e}U$. If $U=\varnothing$ we simply say that $\{A,B\}$ is a $\mathcal{{K}}-pair$.
Being half of $\mathcal{{K}}_{U}$-pair turns out to be a degree-theoretic notion, and the remarkably simple characterization obtained by Kalimullin hints at their usefulness in definability results.
Proposition 6.2 (Kalimullin [24]).
Let $A,B,C,U\in\mathcal{{P}}(\mathbb{{N}})$
  1. (1)
    If $\{A,B\}$ is a $\mathcal{{K}}_{U}$-pair and $C\leq_{e}B$, then $\{A,C\}$ is a $\mathcal{{K}}_{U}$-pair.
  2. (2)
    If $\{A,B\}$ is a $\mathcal{{K}}_{U}$-pair and $\{A,C\}$ is a $\mathcal{{K}}_{U}$-pair, then $\{A,B\oplus C\}$ is a $\mathcal{{K}}_{U}$-pair.
  3. (3)
    $\{A,B\}$ is a $\mathcal{{K}}_{U}$-pair if and only if
    $\forall X\;\deg_{e}(X)=\deg_{e}(A\oplus U\oplus X)\land\deg_{e}(B\oplus U\oplus X).$
Observe that $\mathcal{{K}}_{U}$-pairs are symmetric, so the proposition above holds equally if we interchange the roles of $A$ and $B$. The first item implies that the notion of $\mathcal{{K}}_{U}$-pairs is degree-theoretic, and the first and second combined state that set of $\mathcal{{K}}_{U}$-partners of half of a $\mathcal{{K}}_{U}$-pair forms an ideal in the enumeration degrees. The converse of the third item also holds, which gives the definability of $\mathcal{{K}}$-pairs in the enumeration degrees. Moreover, it shows that a $K_{U}$-pair $\{A,B\}$ is also a $K_{V}$-pair for any $V\geq_{e}U$ and that it is nontrivial if we also have that $V\ngeq_{e}A,B$. If we restrict our attention to nontrivial $\mathcal{{K}}_{U}$-pairs, we can extract more information.
Proposition 6.3 (Kalimullin [24]).
@@
    $A\leq_{e}\overline{B}\oplus U$ and $\overline{A}\leq_{e}B\oplus U^{\diamond}$. In particular, $A\leq_{e}B^{\diamond}\oplus U$ and $A^{\diamond}\leq_{e}B\oplus U^{\diamond}$.
@@
    The set $A\oplus U$ is a strong quasiminimal cover of $U$.
A nontrivial semicomputable pair $\{A,\overline{A}\}$ is not only a $\mathcal{{K}}$-pair, but a maximal $\mathcal{{K}}$-pair [12]: if $C\geq_{e}A$, $D\geq_{e}\overline{A}$, and $\{C,D\}$ is a $\mathcal{{K}}$-pair, then $A\equiv_{e}C$ and $\overline{A}\equiv_{e}D$. Cai et al. [6] proved that these are the only maximal $\mathcal{{K}}$-pairs up to enumeration degree (we define maximal $\mathcal{{K}}$-pairs to always be nontrivial), concluding that total degrees are definable. This can be relativized with a total oracle $T$ to obtain that the degrees of maximal $\mathcal{{K}}_{T}$-pairs are exactly the degrees of nontrivial semicomputable pairs relative to $U$. In general, if $\{A,\overline{A}\}$ is a semicomputable pair and $A,\overline{A}\nleq_{e}U$, we have that $\{A\oplus U,\overline{A}\oplus U\}$ is a maximal $\mathcal{{K}}_{U}-pair$: if $\{C,D\}$ is a (necessarily nontrivial) $\mathcal{{K}}_{U}-pair$ with $A\oplus U\leq_{e}C$ and $\overline{A}\oplus U\leq_{e}D$, by Proposition 6.2.1, $\{A,D\}$ is a nontrivial $\mathcal{{K}}_{U}$-pair and by Proposition 6.3.a we have that $D\leq_{e}\overline{A}\oplus U$. Similarly, $C\leq_{e}A\oplus U$. It is not known if all maximal $\mathcal{{K}}_{U}$-pairs can be obtained in this way when $U$ is not total.
6.2. A Borel Invariant Function that is not Uniformly Invariant
Now we have all the ingredients to construct the $e$-invariant function that is not uniformly invariant. The following lemma provides the combinatorial core of the counterexample: a uniformly $e$-invariant function that maps one half of a $\mathcal{{K}}_{U}$-pair to one of its $\mathcal{{K}}$-partners must be constant on that degree.
@@ 6.2. A Borel Invariant Function that is not Uniformly Invariant
If $\bm{a}\leq\bm{b}$, then $\bm{u}=(\bm{a}\vee\bm{u})\wedge(\bm{b}\vee\bm{u})=\bm{a}\vee\bm{u}$. This means that $\bm{a}\leq\bm{u}$, so $\{A,B\}$ is a trivial $\mathcal{{K}}_{U}$-pair, contrary to our assumption. If instead $\bm{a}^{\diamond}\leq\bm{b}$, the monotonicity of the skip implies that $\bm{u}^{\diamond}\leq\bm{a}^{\diamond}$. Combined with Proposition 6.2.a,
@@
So, to prove that a function is not uniformly $e$-invariant, it is enough to build a function that maps a cofinal set of halves of nontrivial relativized $\mathcal{{K}}$-pair to their $\mathcal{{K}}$-partners, we devote the rest of the section to the construction of a counterexample that works on every cone and to the proof that no Borel function is equivalent to it on a cone. The main obstacle for the construction is finding, on input $A$, if there are sets $B$ and $U$ such that $\{A,B\}$ is a nontrivial $\mathcal{{K}}_{U}$-pair so that we can map $A$ to $B$ while at the same time ensuring that if $C\equiv_{e}A$, we make a consistent choice $D\equiv_{e}B$ as the output on $C$. The second part can be managed by restricting our attention to maximal $\mathcal{{K}}_{U}$-partners, as those inhabit a single degree. The next lemma takes care of the first part; the key idea is to ensure that a unique $U$ will work by only looking for $\mathcal{{K}}_{U}$-pairs $\{A,B\}$ where $U$ has total degree and $U<_{e}A,B$.
@@
There is a Borel $e$-invariant function $F\colon\mathcal{{P}}(\mathbb{{N}})\to\mathcal{{P}}(\mathbb{{N}})$ such that $F$ is not equivalent to any Borel uniformly $e$-invariant function on any cone.
@@
Of course, we need to ensure that it is single-valued and Borel. We will construct the function $F$, on input $A$, according to the following instructions. First, search for a total set $X\oplus\overline{X}$ below $A$ and a candidate $B$ below $A^{\diamond}$ such that $\{A,B\}$ is a maximal $K_{X\oplus\overline{X}}$-pair. If found, map $A$ to $B$. Otherwise, map $A$ to $A$.
While the procedure to obtain $F(A)$ from $A$ is not computable, it is arithmetical. To make things precise, we define $F(A)$ by the following procedure:
@@
A rough calculation using Theorem 2.6(III) in [24] gives a bound of $\Sigma^{0}_{4}(A^{\prime})$ on the complexity of $F(A)$. The uniqueness (up to $e$-degree) of maximal $\mathcal{{K}}$-partners guarantees that the function is $e$-invariant, and step 2 ensures that $F$ is not locally constant. Hence, by Lemma 6.4, that $F$ is not uniformly $e$-invariant.
We need to work a bit harder to show that is not equivalent, on any upper cone, to a Borel uniformly $e$-invariant function. The idea is that, if it were, it would give us a Borel choice function for the enumeration degrees of halves of maximal $\mathcal{{K}}$-pairs. With such a choice function, we can define a wellorder of type $\omega$ on each such degree, which leads us to a contradiction. We proceed with the details:
Suppose towards a contradiction that there are a set $C\subseteq\mathbb{{N}}$ and a Borel uniformly $e$-invariant function $H\colon\mathcal{{P}}(\mathbb{{N}})\to\mathcal{{P}}(\mathbb{{N}})$ such that $H(X)\equiv_{e}F(X)$ for all $X\geq_{e}C$. Fix a Borel code $T$ for $H$ and let
$\operatorname{\mathcal{{K}}Max_{C}}=\{A\mid\exists B\ \exists U\;(C\leq_{e}U\oplus\overline{U}\mbox{ and }\{A,B\}\mbox{ is a maximal }\mathcal{{K}}_{U\oplus\overline{U}}\mbox{-pair}\}.$
For $A\in\operatorname{\mathcal{{K}}Max_{C}}$, we know that $H\restriction_{\deg_{e}(A)}$ is constant by Lemma 6.4. In particular, for all $X\geq_{e}C$, we have that $A\equiv_{e}X$ if and only if $H(H(A))=H(H(X))$. Define $R(X)$ to be the least $n\in\mathbb{{N}}$ such that $X=\Gamma_{n}(H(H(X))$. Observe that $R$ is defined for every set $X\geq_{e}C$ because we always have that $X\equiv_{e}H(H(X))$. Moreover, $R$ induces a wellorder $\leq$ on $\deg_{e}(A)$ of type $\omega$ by $X\leq Y$ if $R(X)\leq R(Y)$.
Denote by $L_{X}$ the left cut induced by the characteristic function of $X$ on the lexicographical ordering of $2^{<\omega}$. Observe that there is a single enumeration operator $\Gamma$ such that $\Gamma(X)=L_{X}$ for all $X\subseteq\mathbb{{N}}$. For a sufficiently $C\oplus T$-generic $G\in 2^{\omega}$, $L_{G},\overline{L_{G}}\nleq_{e}C\oplus\overline{C}$, and there is some $n\in\mathbb{{N}}$ such that $R(L_{G}\oplus C\oplus\overline{C})=n$. So, by genericity, there must be some $k\in\mathbb{{N}}$ such that $G\restriction k\Vdash R(L_{G}\oplus C\oplus\overline{C})=n$. Let $G^{*}\in 2^{\omega}$ be the infinite binary sequence that results from changing the $k+1$ bit of $G$. That is, $G(j)=G^{*}(j)$ if and only if $j\neq k$. Since we only changed one bit, $G^{*}$ is also generic and $G\equiv_{e}G^{*}$. Moreover, $L_{G}\equiv_{e}L_{G^{*}}$. To see this, assume without loss of generality that $G(k)=1$ and $G^{*}(k)=0$, and fix $\sigma=G\restriction k=G^{*}\restriction k$. Then $\tau\in L_{G^{*}}$ if and only if $|\tau|<k+1$ and $\tau<_{lex}\sigma$, or $|\tau|\geq k$ and the string $\tau^{*}\in L_{G}$, where
$\tau^{*}(n)=\begin{cases}\tau(n)&\mbox{if }n<|\tau|\mbox{ and }n\neq k-1.\\ 1&\mbox{if }n=k-1.\end{cases}$
Putting everything together, we have that $L_{G}\oplus C\oplus\overline{C},L_{G^{*}}\oplus C\oplus\overline{C}\in\operatorname{\mathcal{{K}}Max_{C}}$ because $L_{G},L_{G^{*}}$ are semicomputable sets, $H(H(L_{G}\oplus C\oplus\overline{C}))=H(H(L_{G^{*}}\oplus C\oplus\overline{C}))$ because $L_{G}\equiv_{e}L_{G^{*}}$, and $R(L_{G}\oplus C\oplus\overline{C})=R(L_{G^{*}}\oplus C\oplus\overline{C})$ because $G$ and $G^{*}$ are generics with $\sigma\prec G,G^{*}$. However, this is a contradiction as $L_{G}$ and $L_{G^{*}}$ are different sets (one contains $\sigma^{\smallfrown}1$ while the other does not), so one of them is not equal to $\Gamma_{n}(H(H(L_{G}\oplus C\oplus\overline{C})))$.
@@ 7. Further Questions
Is there an $e$-invariant function $f$ such that $f$ is regressive and $f\restriction_{\deg(A)}$ is not constant for all $A\subseteq\omega$?
A different approach to study determinacy in the enumeration degrees was suggested by Slaman (personal communication). The failure of the Cone Theorem in the enumeration degrees leaves open whether some of the consequences of determinacy can still be proven. For example, a classical application of TD shows that there is a cone $\mathcal{{C}}$ of Turing degrees such that every cone $\mathcal{{D}}_{T}(\geq\bm{d})$ above an element $\bm{d}\in\mathcal{{C}}$ has the same first-order theory. The same argument can be used to show that, under TD, there is a total enumeration degree $\bm{t}$ such that for all total degrees $\bm{d}\geq\bm{t}$, the cone of enumeration degrees $\mathcal{{D}}_{e}(\geq\bm{d})$ has the same first-order theory as $\mathcal{{D}}_{e}(\geq\bm{t})$. However, this is false for arbitrary enumeration degrees. For example, the proof of downwards density of $\mathcal{{D}}_{e}$ relativizes to any total degree; thus, for a total degree $\bm{t}$, $\mathcal{{D}}_{e}(\geq\bm{t})$ has no minimal degrees [14]. In contrast, by a theorem of Calhoun and Slaman [7], above every total degree there is a degree $\bm{a}$ such that $\mathcal{{D}}_{e}(\geq\bm{a})$ has a minimal degree.
@@
The same question might be asked about other classes of degrees. For example, the continuous degrees are strictly between the total and the cototal degrees under the subset relation and are first-order definable [1]. Given that all the known proofs that there is a non-total continuous degree are topological in nature, understanding the theory of the cone above a non-total continuous degree could provide a useful technical tool.
@@ Acknowledgments
Thanks to Mariya Soskova for many helpful discussions and advice, to Josiah Jacobsen-Grocott for sharing his insight, to Patrick Lutz for pointing out an omission on the original version of this paper and suggesting an improvement to Theorem 6.6, and to Beth Pcheco for her thoughtful suggestions. The author was supported by the Secretaría de Ciencia, Humanidades, Tecnología e Innovación of México through scholarship number 836694.